OpenStudy (anonymous):

When the number of marbles Jill has is subtracted from 20, the result is the same as adding 16 to the number of marbles. How many marbles does Jill have?

OpenStudy (anonymous):

please help. fan and medal.

OpenStudy (anonymous):

what are your answer choices

OpenStudy (anonymous):

it dosent have any. lol

OpenStudy (anonymous):

i have to write how many marbles she has.

OpenStudy (anonymous):

@jakejjke u there?

OpenStudy (anonymous):

20 - x = x + 16 Solve that we get: 4 = 2x x = 4/2 x = 2 Double-check: 20 - 2 = 18 16 + 2 = 18 Answer: Jill has 2 marbles

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

no problem